OpenStudy (sphott51):

2/(w^2 +2w -24) +5/(w^2 +w -20) = 1/(w^2 +11w +30)

OpenStudy (itz_sid):

I think you factor out the denominators and then then find a common denominator from there.

OpenStudy (sphott51):

and then I find the LCD right?

OpenStudy (sphott51):

I got 2w +10 +5w +30 = w-4 after LCD and crossing out aand distributing.

OpenStudy (robtobey2):

\[\frac{2 (3 w+22)}{(w-4) (w+5) (w+6)}=0 \]3w + 22 = 0 Solve for w.

OpenStudy (sphott51):

-22/3 =w

OpenStudy (sphott51):

2 is 0 right..? or did i do that wrong?